# RoboReflect: Robotic Reflective Reasoning for Grasping Ambiguous-Condition Objects

### v1

@@ Abstract
As robotic technology rapidly develops, robots are being employed in an increasing number of fields. However, due to the complexity of deployment environments or the prevalence of ambiguous-condition objects, the practical application of robotics still faces many challenges, leading to frequent errors. Traditional methods and some LLM-based approaches, although improved, still require substantial human intervention and struggle with autonomous error correction in complex scenarios. In this work, we propose RoboReflect, a novel framework leveraging large vision-language models (LVLMs) to enable self-reflection and autonomous error correction in robotic grasping tasks. RoboReflect allows robots to automatically adjust their strategies based on unsuccessful attempts until successful execution is achieved. The corrected strategies are saved in a memory for future task reference. We evaluate RoboReflect through extensive testing on eight common objects prone to ambiguous conditions of three categories. Our results demonstrate that RoboReflect not only outperforms existing grasp pose estimation methods like AnyGrasp and high-level action planning techniques using GPT-4V but also significantly enhances the robot’s ability to adapt and correct errors independently. These findings underscore the critical importance of autonomous self-reflection in robotic systems while effectively addressing the challenges posed by ambiguous environments.
@@ I INTRODUCTION
With the rapid development of artificial intelligence, embodied AI and robotics technology have made significant strides across various fields. A notable instance is robotic arm grasping tasks, which have become a vital application in industrial automation and service robotics [1, 2, 3]. The recent rise of Large Language Models (LLMs) has further broadened the potential of these technologies. Compared to traditional methods, LLM-based robotic arms can understand tasks from natural language instructions and autonomously generate grasping policies. This capability makes robotic arms more flexible when facing diverse and dynamically changing tasks. Furthermore, this also lays the foundation for robotic arms to grasp complex, long-horizon tasks. Therefore, researchers are now exploring ways to integrate LLMs into robotic systems, envisioning them as the “brains” of embodied intelligence, capable of facilitating actions in the real world.
However, in real-world scenarios, the properties of daily necessities often differ from those in experimental training data (e.g., Empty tissue bags, deformed bottles, and fragile toys). Some objects whose properties change with use undergo corresponding changes at different conditions, consequently resulting in errors in the grasping (e.g., fulfilled tissue bags and empty tissue bags). Other objects cannot be grasped based on prior training due to factors like material composition or edibility (e.g., ice cream with bar). We refer to such objects as ambiguous-condition objects, where both robotic arm systems and LLMs struggle to determine the object’s current condition at the first time, often leading to grasping failures or unreasonable actions. Therefore, whether applying traditional methods or LLM-driven approaches, distinguishing between objects with the same label but differing physical properties remains highly challenging for ensuring proper grasping. How to reduce these misjudgments remains a key problem in more general and practical applications.
@@
  • RoboReflect utilizes the Reflective Reasoning module with LVLM and a Memory module, enabling the framework to resolve the ambiguous grasping scenario.
  • We categorize ambiguous-condition objects into three types, and test RoboReflect on eight common objects. Ours consistently outperforms AnyGrasp [6] and GPT-4V [7] driven grasp method, which highlights the critical role of autonomous reflection-correction in ambiguous grasping scenarios.
@@ III METHODOLOGY
In this section, we first formulate the problem (Sec. III-A) and then introduce RoboReflect, a framework for autonomous reflection and correction of robotic actions in ambiguous grasping scenario. The framework comprises four modules: the visual processing and action modules generate robot actions based on visual observations and instructions (Sec. III-B); the judgment module assesses task completion from environmental feedback (Sec. III-C); the reflective reasoning module analyzes failures using LVLM (Sec. III-D); and the memory module records successful strategies to prevent future errors (Sec. III-E). The framework is shown in Fig. 1.
@@ III-A Problem Statement
Specifically, first, by applying the RGB-Depth sensors on the robotic arm, we captured a set of RGB-D images $\bm{ID_{N}}$ (Since processing image sequence data is not involved in our pipeline, unless otherwise specified, a single character will be used to represent the sequence data in this paper, i.e., $\bm{ID}$ will represent $\bm{ID_{N}}$). We then use the Visual Processing module ($\bm{VP}$) to process the RGB-D data into RGB data $\bm{I}$ and 3D spatial position $\bm{S}$, which LVLM can understand. Simultaneously, an instruction text prompt $\bm{Ins}$ is added, forming the pre-task information $[\bm{I},\bm{S},\bm{Ins}]$ as the input data for $\bm{AM}$. Besides the action output, we also store the action sequence’s corresponding action images, represented as $\bm{I_{act}}$. These action images are then input into the final $\bm{JM}$ to generate the binary grasp determination $[\bm{G_{S}},\bm{G_{P}}]$. The whole process can be presented as:
$[\bm{G_{S}},\bm{G_{P}}]=\bm{JM}[\bm{AM}[\bm{VP}(\bm{I},\bm{S}),\bm{Ins}],\bm{S},\bm{Ins}]$ (2)
Once we obtain the binary grasp determination, if $\bm{G_{S}}\cup\bm{G_{P}}=1$, it indicates that the object has been successfully grasped, and we store the prompt in the memory for future retrieval and reference. However, if $\bm{G_{S}}\cup\bm{G_{P}}=0$, further adjustments are needed through the $\bm{RRM}$ until the $\bm{JM}$ determines that $\bm{G_{S}}\cup\bm{G_{P}}=1$, storing the object description $O$ and finishing the grasp.
@@ III-B Converting RGB-D images & Action Planning
Visual processing module ($\bm{VP}$). To enable an embodied robot to interact effectively with the environment and plan appropriate actions, it is crucial to help the robot understand the spatial relationships of objects. Initially, the LVLM $\mathcal{M}$ utilizes visual observations $\bm{I}$ to identify objects $\bm{O}$ within the scene and generate their captions along with 2D bounding boxes. Subsequently, we use the detected bounding boxes to prompt the SAM[20] to derive segmentation masks, thereby refining the spatial occupancy. This process enables us to extract the point cloud of each object by leveraging the camera parameters and depth information. Consequently, we can obtain the 3D location and bounding boxes of each object, collectively referred to as the 3D spatial position $\bm{S}$.
Action module ($\bm{AM}$) is crucial for converting the data information to the robot’s arm actions. We empower embodied robots with a series of actions, each primitive action is in the form of a predefined API, such as “Move”, “GraspOn” and “GraspOff”. By combining these atomic actions, the robot can accomplish some basic tasks. Specifically, we prompt LVLM $\mathcal{M}$ to decompose task instructions $\bm{Ins}$ into step-by-step instructions based on input scene information $[\bm{I},\bm{S}]$, each of which specifies a predefined primitive action. Besides action instructions for the arm, this module also outputs the action’s corresponding images which are captured by the RGB camera on the robot arm as the action images $\bm{I_{act}}$.
@@ III-C Status Determining by Judgment Module
During the action process, $\bm{I_{act}}$ are stored to present the state of the environment following each action. In the $\bm{JM}$, a series of action images $\bm{I_{act}}$, along with the instruction text prompt $\bm{Ins}$ and 3D position $\bm{S}$, are encapsulated in a prompt to provide the LVLM $\mathcal{M}$ with the necessary information to infer the task finished status. The $\bm{JM}$ then evaluates the grasping success status based on two predefined criteria $[\bm{G_{S}},\bm{G_{P}}]$. When dealing with objects under ambiguous conditions, the judgment primarily hinges on the following aspects: 1) Unsuccessful grasping often stems from changes in the object’s properties during manipulation, such as deformations or structural separations, which ultimately lead to failure. 2) Incorrect grasping positions, while possibly effective for manipulation, are generally considered inappropriate by human experience. This includes grasping edible parts or areas that may pose safety risks, making such positions undesirable from a functional and safety perspective. Therefore, along with $\bm{I_{act}}$ and $\bm{Ins}$ into the LVLM, two additional textual questions are input into the LVLM module: 1) Was the robotic arm’s grasp successful? 2) Does the grasping position align with human experience?
When the $\bm{JM}$ determines that the task is unsuccessful, i.e. $\bm{G_{S}}\cup\bm{G_{P}}=0$, it forwards the action images $\bm{I_{act}}$ and requests the Reflective Reasoning Module to initiate reflection (c.f. Sec. III-D). This core component of the framework represents the first step toward autonomous reflection and reasoning. If the evaluation is deemed successful, i.e. $\bm{G_{S}}\cup\bm{G_{P}}=1$, by the $\bm{JM}$, the object $\bm{O_{i}}$ information will be stored within the memory module (c.f. Sec. III-E) for generalizing to grasp other objects in the future.
@@ III-D Correction by Reflective Reasoning Module
Self-Reflective Module. When the Judgement Module determines a grasping failure, analyzing the cause of the failure becomes the key focus which will assist the robot to correct the action. For objects in ambiguous conditions where the LVLM $\mathcal{M}$ cannot accurately determine their properties, the robotic arm often struggles to execute the correct grasping action. In such cases, a lot of key information must be inferred through interaction with the object and observation of feedback in order to perform the correct grasp. Therefore, we introduce a self-reflective Module to analyze and discuss the specific causes of grasping failures, wrapping them into prompts understandable by the LVLM $\mathcal{M}$. The inputs to the self-reflective module include not only the environment feedback $\bm{I_{act}}$ but also the task’s input object’s text description $\bm{O}$ and text instructions $\bm{Ins}$ as prior information.
Figure 2: Reflective Reasoning Module. The Left part is the Self-Reflective Module and the right part is the Discussion Module. (Zoom in to view clearly.)
@@
As shown on the left of Fig. 2, the LVLM $\mathcal{M}$ first analyzes the object description $O$ to understand its intrinsic properties. Next, $\mathcal{M}$ considers the object’s context within the environment, linking potential states to the action images $\bm{I_{act}}$. Given that ambiguous objects often pose two main challenges—changes in object properties and sub-optimal grasping $\mathcal{M}$ is guided to reason in these two directions when determining the error cause $\bm{Y}$. Once the LVLM $\mathcal{M}$ identifies the error cause and proposes action corrections $\bm{P}$ through reasoning, these suggestions must comply with the inferred constraints of the object. This process yields the final reflection result $\bm{R}$ for the robot.
@@
To enable RoboReflect to robustly handle objects in various ambiguous conditions, we draw inspiration from recent efforts in peer-rating by LLMs[23], and introduce the Discussion Module with another LVLM $\mathcal{M}_{D}$ to supervise the results of the Self-Reflective Module. More specifically, after completing the introspective reasoning process, instead of immediately planning new actions, we utilize a newly introduced large-language model, to evaluate the result $R$. This discussion is formalized as $\bm{R^{\prime}}=\mathcal{M}_{D}(\bm{R},\bm{I_{act},\bm{Ins}})$, where $\bm{R^{\prime}}$ encapsulates the discussion outcome.
The discussion process primarily involves two steps,, as shown in the right of Figure 2: The first step is to evaluate the correctness of $\bm{R}$ by the discussion LVLM $\mathcal{M}_{D}$ with multi-turns Q&A. If $\bm{R}$ is determined to be correct by the LVLM $\mathcal{M}$, the reflection result will be retained. However, if any issues are identified, we proceed to update and improve $\bm{R}$. Essentially, this approach is premised on a re-analysis of the causes of errors based on the results of reflection and reasoning. It serves to enhance our understanding of objects under ambiguous conditions, thereby avoiding any oversight of grasping constraints. The reflect suggestion $\bm{R^{\prime}}$, is fed back to the LVLM $\mathcal{M}$ for next time grasp trying. It serves as an explanation for the errors encountered and an understanding of objects under ambiguous conditions, thereby assisting the model in planning accurate grasping actions.
@@ III-E Memory Module
Successful task execution signifies that the system has mastered the operable attributes of the object. To avoid wasteful repetition of reflection and to facilitate efficient and accurate task completion in the future, it is crucial to retain the results of reflection for subsequent use. Therefore, we employ a memory module to record the details of successfully executed tasks. This will serve as action experience before assisting future task execution. The memory is structured as the dictionary data structure, where each key represents an object’s description $\bm{O}$ with which there has been interaction, and the corresponding value encapsulates the understanding $R$ derived from interacting with the object. Given that the same object may present different ambiguous conditions in different scenarios, to prevent memory deception leading to errors, each memory is only associated with the actions of that stage. If the scenario changes, all memories are cleared.
@@ IV EXPERIMENTS
Figure 3: Objects descriptions. The upper half of tissue bag is often empty due to the exhaustion of materials, which makes them susceptible to deformation. A closed-lid cup refers to a standard cup with a securely fastened lid. In the case of an open-lid cup, the lid is not tightly screwed on, which can lead to separation between the lid and the body of the cup. Cookies are fragile and prone to breakage. The upper half of a hard drive is marked with signs indicating that it should not be touched. A sealed cup noodles represents the standard configuration of instant noodle cups. An unsealed cup noodles has an unsealed top, making it vulnerable to spillage and deformation. The edible part of an ice cream bar should not be touched.
Figure 4: The comparison between failed and successful grasping cases. We present the grasping results of six objects before and after using our RoboReflect model. Before reflection, the grasping attempts often failed due to an inability to understand the object’s properties, resulting in either a failure to grasp or an incorrect grasp position. After reflection, the model was able to successfully grasp the objects.
@@ IV-A Experimental Setup
Metrics. To quantify our performance, we use the success rate of grasping ambiguous-condition objects for evaluation. The success rate not only considers the grasping state $\bm{G_{S}}$ but also the grasping position $\bm{G_{P}}$ correct or not. The correct grasping position $\bm{G_{P}}$ of objects is always defined by the human experience. Based on the inherent human experience of the ambiguous condition grasping, we further provide the Human-experience Evaluation Criteria (HEC). We categorize objects prone to ambiguous conditions into three types, the details descriptions are shown in Fig. 3:

### v2

# RoboReflect: A Robotic Reflective Reasoning Framework for Grasping Ambiguous-Condition Objects

@@ Abstract
As robotic technology rapidly develops, robots are being employed in an increasing number of fields. However, due to the complexity of deployment environments or the prevalence of ambiguous-condition objects, the practical application of robotics still faces many challenges, leading to frequent errors. Traditional methods and some LLM-based approaches, although improved, still require substantial human intervention and struggle with autonomous error correction in complex scenarios. In this work, we propose RoboReflect, a novel framework leveraging large vision-language models (LVLMs) to enable self-reflection and autonomous error correction in robotic grasping tasks. RoboReflect allows robots to automatically adjust their strategies based on unsuccessful attempts until successful execution is achieved. The corrected strategies are saved in the memory for future task reference. We evaluate RoboReflect through extensive testing on eight common objects prone to ambiguous conditions of three categories. Our results demonstrate that RoboReflect not only outperforms existing grasp pose estimation methods like AnyGrasp and high-level action planning techniques ReKep with GPT-4V but also significantly enhances the robot’s capability to adapt and correct errors independently. These findings underscore the critical importance of autonomous self-reflection in robotic systems while effectively addressing the challenges posed by ambiguous-condition environments.
@@ I INTRODUCTION
With the rapid development of artificial intelligence, embodied AI and robotics technology have made significant strides across various fields. A notable instance is robotic arm grasping tasks, which have become a vital application in industrial automation and service robotics [1, 2, 3]. The recent rise of Large Language Models (LLMs) has further broadened the potential of these technologies. Compared to traditional methods, LLM-based robotic arms can understand tasks from natural language instructions and autonomously generate grasping policies. This capability makes robotic arms more flexible when facing diverse and dynamically changing tasks. Furthermore, this also lays the foundation for robotic arms to deal with complex, long-horizon manipulation tasks. Therefore, researchers are now exploring ways to integrate LLMs into robotic systems, envisioning them as the “brain” of embodied intelligence, capable of facilitating actions in the real world.
In real-world scenarios, the properties of daily necessities often differ from those in experimental training data (e.g., Empty tissue bags, deformed bottles, and fragile toys). Some objects whose properties change with use undergo corresponding changes at different conditions, consequently resulting in errors in the grasping (e.g., fulfilled tissue bags and empty tissue bags). Additionally, achieving reliable grasping performance necessitates deeper understanding of objects due to factors such as material composition or edibility (e.g., ice cream on a stick).
We refer to such objects as ambiguous-condition objects, where both robotic arm systems and LLMs struggle to determine the object’s current condition to calculate the affordance point or grasp pose at the first time, often leading to grasping failures or unreasonable actions. Therefore, whether applying traditional methods or LLM-driven approaches, distinguishing between objects with the same label but differing physical properties remains highly challenging for ensuring proper grasping. How to reduce these misjudgments remains a key problem in more general and practical applications.
@@
  • RoboReflect utilizes the Reflective Reasoning module with LVLM and a Memory module, enabling the framework to resolve the ambiguous grasping scenario and incorporates a memory mechanism to handle ambiguous-condition issues that have already been encountered.
  • We categorize ambiguous-condition objects into three types, and test RoboReflect on eight common objects. Ours consistently outperforms ReKep [6], AnyGrasp [7] and GPT-4V [8] driven grasp method, which highlights the critical role of autonomous reflection-correction in ambiguous grasping scenarios.
@@ III METHODOLOGY
In this section, we first formulate the problem (Sec. III-A) and then introduce RoboReflect, a framework for autonomous reflection and correction of robotic actions in the ambiguous grasping scenario. The framework comprises four modules: the visual processing and action modules generate robot actions based on visual observations and instructions (Sec. III-B); the judgment module assesses task completion from environmental feedback (Sec. III-C); the reflective reasoning module analyzes failures using LVLM (Sec. III-D); and the memory module records successful strategies to prevent future errors (Sec. III-E). The framework is shown in Fig. 1.
@@ III-B Converting RGB-D images & Action Planning
Visual processing module ($\bm{VP}$). To establish object-centric spatial representations for embodied interaction, our framework hierarchically processes visual observations through three sequential transformations. The process begins with the vision-language model $\mathcal{M}$ analyzing RGBD inputs $\bm{I}\in\mathbb{R}^{H\times W\times 4}$ to detect $N$ objects $\mathcal{O}=\{o_{i}\}_{i=1}^{N}$, generating both semantic captions $\bm{C}$ and preliminary 2D bounding boxes $\bm{B}_{\text{2D}}\in\mathbb{R}^{N\times 4}$ through its frozen detection head. These 2D coordinates then serve as geometric prompts for SAM[21], which produces pixel-precise instance masks that refine object boundaries beyond the initial detection. By back-projecting the masked depth data via camera intrinsics, we get the 3D position and bounding boxes of each object, collectively referred to as the 3D spatial position $\bm{S}$.
Action module ($\bm{AM}$) serves as the core component of the robotic execution system, responsible for translating high-level instructions into executable motion sequences for robotic manipulators. In our embodied agent implementation, we systematically encapsulated a set of atomic actions (e.g., “Move”, “GraspOn” and “GraspOff”) through predefined API libraries, establishing fundamental units for task decomposition and recomposition. The system parses natural language instructions $\bm{Ins}$ by analyzing scene information $[\bm{I},\bm{S}]$. For instance, the command “pick up the water cup” would be decomposed into an atomic action sequence: Move(’cup’), GraspOn() and Move(). Notably, during the execution of the action, $\bm{AM}$ records the visual data collected by the RGB camera at the end of the robot arm. These action execution image sequences $\bm{I_{act}}$ provide the system with visual evidence for action verification.
@@ III-C Status Determining by Judgment Module
In the action process, the images $\bm{I_{act}}$ is stored to present the state of the environment action. In the $\bm{JM}$, the last frame image $\bm{I_{act}}$, along with the instruction text prompt $\bm{Ins}$ and 3D bounding box $\bm{S}$, are encapsulated in a prompt to provide the LVLM $\mathcal{M}$ with the necessary information to infer the task finished status. The $\bm{JM}$ then evaluates the grasping success status based on two predefined criteria $[\bm{G_{S}},\bm{G_{P}}]$. When dealing with objects under ambiguous conditions, the judgment primarily hinges on the following aspects: 1) Unsuccessful grasping often stems from changes in the object’s properties during manipulation, such as deformations or structural separations, which ultimately lead to failure. 2) Incorrect grasping positions, while possibly effective for manipulation, are generally considered inappropriate by human experience. This includes grasping edible parts or areas that may pose safety risks, making such positions undesirable from a functional and safety perspective. Therefore, along with $\bm{I_{act}}$ and $\bm{Ins}$ into the LVLM, two additional textual questions are input into the LVLM module: 1) Was the robotic arm’s grasp successful? 2) Does the grasping position align with human experience?
When the $\bm{JM}$ determines that the task is unsuccessful, i.e. $\bm{G_{S}}\cup\bm{G_{P}}=0$, it forwards the action images $\bm{I_{act}}$, $\bm{Ins}$ to requests the Reflective Reasoning Module to initiate reflection (c.f. Sec. III-D). This core component of the framework represents the first step toward autonomous reflection and reasoning. If the evaluation is deemed successful, i.e. $\bm{G_{S}}\cup\bm{G_{P}}=1$, by the $\bm{JM}$, the object $\bm{O_{i}}$ description information will be stored within the memory module (c.f. Sec. III-E) for generalizing to grasp other objects in the future task.
@@ III-D Correction by Reflective Reasoning Module
Self-Reflective Module. When the judgment module determines the grasping failure, analyzing the cause of the failure becomes the key focus which will assist the robot to correct the action. For objects in ambiguous conditions where the LVLM $\mathcal{M}$ cannot accurately determine their properties, the robotic arm often struggles to execute the correct grasping action. In such cases, a lot of key information must be inferred through interaction with the object and observation of feedback in order to perform the correct grasp pose. Therefore, we introduce a self-reflective module to analyze and discuss the specific causes of grasping failures, wrapping them into prompts understandable by the LVLM $\mathcal{M}$. The inputs to the self-reflective module include not only the environment feedback $\bm{I_{act}}$ but also the task’s input object’s description $\bm{O}$ and text instructions $\bm{Ins}$ as prior information.
@@
As shown in Fig. 1, the LVLM $\mathcal{M}$ first analyzes the object description $O$ to understand objects intrinsic properties. Next, $\mathcal{M}$ considers the object’s context within the environment, linking potential states to the action images $\bm{I_{act}}$. Given that ambiguous objects often pose two main challenges—changes in object properties and sub-optimal grasping $\mathcal{M}$ is guided to reason in these two directions when determining the error cause $\bm{Y}$. When the LVLM identifies the cause of the error and constrains it through the intrinsic properties of the object description, the model returns the action corrections description $\bm{P}$. At the same time, we combine the identified error description $\bm{Y}$ and $\bm{P}$ to form $\bm{R}=\{\bm{Y},\bm{P}\}$, which is then output as the result.
Figure 2: Objects descriptions. The upper half of tissue bag is often empty due to the exhaustion of materials, which makes them susceptible to deformation. A closed-lid cup has a securely fastened lid. An open-lid cup risks lid-body separation due to loose closure. Cookies are fragile and breakable. The upper half of a hard drive is labeled Untouchable. Sealed cup noodles denote standard intact packaging. Unsealed cup noodles have vulnerable tops prone to spills/deformation. The edible portion of an ice cream bar must remain untouched.
@@
To enable RoboReflect to robustly handle objects in various ambiguous conditions, we draw inspiration from recent efforts in peer-rating by LLMs[24], and introduce the Discussion Module with another LVLM $\mathcal{M}_{D}$ to supervise the results of the Self-Reflective Module. More specifically, after completing the introspective reasoning process, instead of immediately planning new actions, we utilize a module to evaluate the result $R$. This discussion is formalized as $\bm{R^{\prime}}=\mathcal{M}_{D}(\bm{R},\bm{I_{act},\bm{Ins}})$, where $\bm{R^{\prime}}$ encapsulates the discussion outcome.
The discussion process primarily involves two steps,, as shown in the Figure 1: The first step is to evaluate the correctness of $\bm{R}$ by the discussion LVLM $\mathcal{M}_{D}$ with multi-turns Q&A. If $\bm{R}$ is determined to be correct by the LVLM $\mathcal{M}$, the reflection result will be retained. However, if any issues are identified, we will proceed to update and improve $\bm{R}$. Essentially, this approach is premised on a re-analysis of the causes of errors based on the results of reflection and reasoning. It serves to enhance our understanding of objects under ambiguous conditions, thereby avoiding any oversight of grasping constraints. The reflect suggestion $\bm{R^{\prime}}$, is fed back to the LVLM $\mathcal{M}$ for next time grasp trying. It serves as an explanation for the errors encountered and an understanding of objects under ambiguous conditions, thereby assisting the model in planning accurate grasping actions.
@@ III-E Memory Module
Successful task execution signifies that the system has mastered the operable attributes of the object. To avoid wasteful repetition of reflection and to facilitate efficient and accurate task completion in the future, it is crucial to retain the results of reflection for subsequent use. Therefore, we employ a memory module to record the details of successfully executed tasks. This will serve as action experience before assisting future task execution.
The memory is structured as the ‘dictionary’ data structure, where each ‘key’ represents an object’s description $\bm{O}$ with which there has been interaction, and the corresponding ‘value’ encapsulates the understanding $R$ derived from interacting with the object. Given that the same object may present different ambiguous conditions in different scenarios, to prevent memory deception leading to errors, each memory is only associated with the actions of that stage. If the scenario changes, all memories are cleared.
@@ IV EXPERIMENTS
Figure 3: Visual comparison of grasping posture. The green ones are our method, the blue poses are the result of ReKep, and the red ones are the result of AnyGrasp.
Figure 4: The comparison between failed (without reflect) and successful (with reflect) grasping cases. We present the grasping results of six objects before and after using our RoboReflect model. Before reflection, the grasping attempts often failed due to an inability to understand the object’s properties, resulting in either a failure to grasp or an incorrect grasp position. After reflection, the model was able to successfully grasp the objects.
@@ IV-A Experimental Setup
Metrics. To quantify our performance, we use the success rate of grasping ambiguous-condition objects for evaluation. The success rate not only considers the grasping state $\bm{G_{S}}$ but also the grasping position $\bm{G_{P}}$ correct or not. The correct grasping position $\bm{G_{P}}$ of objects is always defined by the human experience. Based on the inherent human experience of the ambiguous condition grasping, we categorize objects prone to ambiguous conditions into three types, the details descriptions are shown in Fig. 2: